# Fully non-linear elliptic equations on noncompact complex manifolds

### v1

@@ Abstract.
In this paper, we obtain the a priori estimates and the existence results for solutions of a general class of fully non-linear equations on noncompact Kähler and Hermitian manifolds. As geometric applications, we can construct Kähler metrics with some prescribed volume forms on strictly pseudoconvex domains; and we can construct complete Chern-Einstein metrics on Hermitian manifolds with bounded geometry.
@@ 1. Introduction
Let $M$ be a complete Kähler manifold of dimension $n$. Let $\alpha$ be a Hermitian metric with bounded geometry (which does not need to be Kähler) on $M$. Fix a real $(1,1)$-form $\chi$ on $(M,\alpha)$, for any $C^{2}$ function $u:\Omega\rightarrow\mathbb{R}$ we have a new real $(1,1)$-form $g=\chi+\sqrt{-1}\partial\bar{\partial}u$, and we can define the endomorphism of $T^{1,0}M$ given by $A_{j}^{i}=\alpha^{i\bar{p}}g_{j\bar{p}}$. In this paper, we consider the equation for $u$ as follows:
@@
is a smooth symmetric function of the eigenvalues of $A$.
@@
First, we aim to obtain a priori $C^{0}$-estimates for (1.1). As we know, any Kähler form $\omega$ can be locally expressed by $\omega=\sqrt{-1}\partial\bar{\partial}\varphi>0$, where $\varphi$ is called the Kähler potential. The key idea of $C^{0}$-estimates is to assume that the Kähler manifold $M$ admits a bounded global Kähler potential, see the following definition:
@@
By generalized maximum principle, we obtain the a priori estimates up to $C^{k+\beta}$:
@@
Let $(M,\alpha)$ be a complete Hermitian manifold with bounded geometry of order k1k-1, $k\geq 4$. Fix a real $(1,1)$-form $\chi$ with bounded geometry of order k1k-1 on $(M,\alpha)$, given $h\in\tilde{C}^{k-2+\beta}(M)$, let $u\in\tilde{C}^{k+\beta}(M)$ satisfy that
@@
Let $\Omega$ be a complete Kähler manifold with a bounded global Kähler potential $\varphi<0$ and $\lim_{x\rightarrow\infty}\varphi(x)=0$. Let $\alpha$ be a Hermitian metric on $\Omega$ with bounded geometry of order k1k-1, where $k\geq 4$. Fix a real $(1,1)$-form $\chi$ with bounded geometry of order k1k-1, and $\chi$ is uniformly equivalent to $\alpha$ on $(\Omega,\alpha)$, given $h\in\tilde{C}^{k-2+\beta}(\Omega\times\mathbb{R})$, where $\beta\in(0,1)$. Consider the following equation on $\Omega$:
@@
Suppose that the above equation satisfies conditions (a1)-(a5) and $F(\alpha^{i\bar{p}}\chi_{j\bar{p}})=0$, assume that $|h(x,0)|=O(\underset{\min}{\lambda}(\alpha^{a\bar{l}}\varphi_{b\bar{l}}))$, and for all $i=1,\ldots,n$
$\textbf{\emph{e}}_{i}$ denotes the $i$th standard basis vector. Then the equation (6.30) exists a solution $u\in\tilde{C}^{k+\beta}(\Omega)$ such that $\chi+\sqrt{-1}\partial\bar{\partial}u$ is uniformly equivalent to $\alpha$.
If, in addition, assuming that $h_{u}(x,u)$ is strictly positive, we can generalize Theorem 1 to Hermitian manifolds:
@@
Suppose that the above equation satisfies conditions (a1)-(a5) and $F(\alpha^{i\bar{p}}\chi_{j\bar{p}})=0$, and for all $i=1,\ldots,n$
$\textbf{\emph{e}}_{i}$ denotes the $i$th standard basis vector. Then the equation (6.33) exists a solution $u\in\tilde{C}^{k+\beta}(\Omega)$ such that $\chi+\sqrt{-1}\partial\bar{\partial}u$ is uniformly equivalent to $\alpha$.
@@
Corollary 1.6.
Let $\Omega$ be a strictly pseudoconvex domain in a complete Kähler manifold $(M,h_{i\bar{j}})$ as in Proposition 3.5 with $l\geq 3$. Define $g=-\log(-\varphi)$ and $\alpha_{i\bar{j}}=g_{i\bar{j}}+\hat{h}_{i\bar{j}}$, where $\varphi$ is the defining function of $\Omega$. Given $h\in\tilde{C}^{l-1+\beta}(\Omega)$, where $\beta\in(0,1)$ is some uniform constant. Consider the following equation on $\Omega$:
(1.12) $\det(\alpha_{i\bar{j}}+u_{i\bar{j}})=e^{h}\det(\alpha_{i\bar{j}})$
@@
The rest of this paper is organized as follows. In Section 2, we provide some preliminaries which may be used later. In Section 3, we construct some complete metrics with bounded geometry. In Section 4, we establish the $C^{0}$ estimates. In Section 5, we establish the a priori estimates up to $C^{2}$. In Section 6, we solve equation (1.1). In Section 7, we discuss the complex Monge-Ampère equation and some geometric applications.
@@ 2. Preliminaries
In this section, we provide some basic results and preliminaries which may be used in the following sections.
Let $(M,g)$ be a Hermitian manifold. Then Chern connection of $g$ is defined as follows: In local holomorphic coordinates $z^{i}$, for a $T^{1,0}$ vector field $X=X^{i}\partial_{i}$ and a $T^{0,1}$ vector field $Y=Y^{\bar{i}}\partial_{\bar{i}}$, where $\partial_{i}:=\frac{\partial}{\partial z^{i}}$, $\partial_{\bar{i}}:=\frac{\partial}{\partial\bar{z}^{i}}$,
$\nabla_{k}X^{i}=\partial_{k}X^{i}+\Gamma_{jk}^{i}X^{j},\;\;\;\nabla_{\bar{k}}X^{i}=\partial_{\bar{k}}X^{i}$
$\nabla_{k}Y^{\bar{i}}=\partial_{k}Y^{\bar{i}},\;\;\;\nabla_{\bar{k}}Y^{\bar{i}}=\partial_{\bar{k}}Y^{\bar{i}}+\overline{\Gamma_{jk}^{i}}Y^{\bar{j}}$
For a $(1,0)$ form $a=a_{i}{\rm d}z^{i}$ and a $(0,1)$ form $b_{\bar{i}}{\rm d}\bar{z^{i}}$,
$\nabla_{k}a_{i}=\partial_{k}a_{i}-\Gamma_{ik}^{j}a_{j},\;\;\;\nabla_{\bar{k}}a_{i}=\partial_{\bar{k}}a_{i}$
$\nabla_{k}b_{\bar{i}}=\partial_{k}b_{\bar{i}},\;\;\;\nabla_{\bar{k}}b_{\bar{i}}=\partial_{\bar{k}}b_{\bar{i}}-\overline{\Gamma^{j}_{ik}}b_{\bar{j}}$
Here $\nabla_{i}:=\nabla_{\partial_{i}}$, and $\Gamma_{jk}^{i}$ are Christoffel symbols, with
$\Gamma_{jk}^{i}=g^{i\bar{l}}\partial_{j}g_{k\bar{l}},$
where $g^{i\bar{l}}g_{j\bar{l}}=\delta_{ij}$,
@@
The tensor operations in this paper follow the Einstein summation convention.
We extend covariant derivatives to act naturally on any type of tensor. For example, if $W$ is a tensor with components $W_{k}^{i\bar{j}}$, then define
$\nabla_{m}W_{k}^{i\bar{j}}=\partial_{m}W_{k}^{i\bar{j}}+\Gamma_{lm}^{i}W_{k}^{l\bar{j}}-\Gamma_{km}^{l}W_{l}^{i\bar{j}}$
$\nabla_{\bar{m}}W_{k}^{i\bar{j}}=\partial_{\bar{m}}W_{k}^{i\bar{j}}+\overline{\Gamma_{lm}^{j}}W_{k}^{i\bar{l}}$
The Hermitian metric $g$ defines a pointwise norm $|\cdot|_{g}$ on any tensor. For example, with $X,Y,a,b$ as above, we define
$|X|_{g}^{2}=g_{i\bar{j}}X^{i}\overline{X^{j}},\;\;|Y|_{g}^{2}=g_{i\bar{j}}Y^{\bar{j}}\overline{Y^{\bar{i}}},\;\;|a|_{g}^{2}=g^{i\bar{j}}a_{i}\overline{a_{j}},\;\;|b|_{g}^{2}=g^{i\bar{j}}b_{\bar{j}}\overline{b_{\bar{i}}}.$
This is extended to any type of tensor. For example, with $W$ as above, we define
$|W|^{2}_{g}=g^{k\bar{l}}g_{i\bar{j}}g_{p\bar{q}}W_{k}^{i\bar{q}}\overline{W_{l}^{j\bar{p}}}.$
Noted that the Chern connection is a connection such that $\nabla g=\nabla J=0$. The associated Chern curvature tensor for the metric $g$ is then defined as
$R_{i\bar{j}k}^{\;\;\;\;\;l}=-\partial_{\bar{j}}\Gamma_{ik}^{l}.$
The Chern-Ricci curvature is defined by
$R_{i\bar{j}}=g^{k\bar{l}}R_{i\bar{j}k\bar{l}}=-\partial_{i}\partial_{\bar{j}}\log\det g.$
Note that if $g$ is not Kähler, then $R_{i\bar{j}}$ may not equal to $g^{k\bar{l}}R_{k\bar{l}i\bar{j}}$.
For two real $(1,1)$-forms $\chi$ and $\chi^{\prime}$ on $(M,g)$, we say that $\chi$ is uniformly equivalent to $\chi^{\prime}$ if and only if
$c^{-1}\chi^{\prime}<\chi<c\chi^{\prime}$
for some constant $c>0$.
Suppose that $F$ is a function defined by
$F(B)=f(\lambda(B))$
for a Hermitian matrix $B=\{B_{i\bar{j}}\}$ with $\lambda(B)\in\Gamma$, throughout the paper we shall use the notation
$F^{ij}(B)=\frac{\partial F}{\partial B_{i\bar{j}}},\;\;\;\;\;F^{ij,kl}(B)=\frac{\partial^{2}F}{\partial B_{i\bar{j}}\partial B_{k\bar{l}}}.$
We denote the minimum eigenvalue of $B$ by $\underset{\min}{\lambda}(B)$.

### v2

@@ Abstract.
In this paper, we establish a priori estimates and existence results for solutions of a general class of fully non-linear equations on noncompact Kähler and Hermitian manifolds. As geometric applications, we construct complete Kähler metrics with prescribed volume forms on strictly pseudoconvex domains, as well as find Einstein metrics on complete noncompact Kähler manifolds and Hessian manifolds with negative first Chern class.
@@ 1. Introduction
Let $(M,\alpha)$ be a complete Hermitian manifold of dimension $n$. Fix a real $(1,1)$-form $\chi$ on $(M,\alpha)$, for any $C^{2}$ function $u:M\rightarrow\mathbb{R}$ we have a new real $(1,1)$-form $g=\chi+\sqrt{-1}\partial\bar{\partial}u$, and we can define the endomorphism of $T^{1,0}M$ given by $A_{j}^{i}=\alpha^{i\bar{p}}g_{j\bar{p}}$. In this paper, we consider the equation for $u$ as follows:
@@
is a smooth symmetric function of the eigenvalues of $A$. Such equations have been studied by Székelyhidi [25] on compact Hermitian manifolds.
@@
Here are some familiar examples of such non-linear operators $f(\lambda)$: When we take $f(\lambda_{1},\ldots,\lambda_{n})=\log\lambda_{1}\ldots\lambda_{n}$, $\Gamma=\Gamma_{n}$, (1.1) is the complex Monge-Ampère equation, Yau [36] first solved such equations on compact Kähler manifolds. Guo-Phong [14] established sharp $L^{\infty}$-estimates of complex Monge-Ampère equations by PDE methods and without pluripotential theory. A related equation, the (n1)(n-1) Monge-Ampère equation was introduced by Fu-Wang-Wu [9], in this case we take $f=\log\widetilde{\lambda_{1}}\ldots\widetilde{\lambda_{n}}$, $\Gamma=\Gamma_{n}$, where
$\widetilde{\lambda_{k}}=\frac{1}{n-1}\sum_{i\neq k}\lambda_{i},$
for each $k$. Note that a class of (n1)(n-1) Monge-Ampère equations falls within our framework (see Section 7); further studies on such equations can be found in [10, 18, 31, 32].
The more general examples are given by $f=\sigma_{k}^{\frac{1}{k}}$ and $f=(\sigma_{k}/\sigma_{l})^{\frac{1}{k-l}}$, $1\leq l<k\leq n$ defined on the cone
$\Gamma_{k}=\{\lambda\in\mathbb{R}^{n}:\sigma_{j}(\lambda)>0,j=1,\ldots,k\},$
where $\sigma_{k}$ is the $k$-th elementary symmetric polynomial
$\sigma_{k}(\lambda)=\sum_{i_{1}<\cdots<i_{k}}\lambda_{i_{1}}\cdots\lambda_{i_{k}},\;\;\;1\leq k\leq n.$
Consider $f=(\sigma_{n}/\sigma_{l})^{\frac{1}{n-l}}(\lambda)=h$, when l=n1l=n-1 and $h$ is constant, Song-Weinkove [28] showed that a solution exists on Kähler manifolds if there is a $\mathcal{C}$-subsolution (see Definition 1.3) for (1.1). Fang-Lai-Ma [11] generalized this result to general $k$.
Since we only require $h_{u}(x,u)\geq 0$ in condition (a5), the $C^{0}$-estimates for (1.1) is a difficulty of the paper. We begin by reviewing some related works. Yau [36] established the $C^{0}$-estimates for Monge-Ampère equations on compact Kähler manifolds and solved the Calabi conjecture. Tosatti-Weinkove [30] used Morse iteration to derive the $L^{\infty}$-estimates for Monge-Ampère equations on compact Hermitian manifolds, generalizing the Calabi conjecture to Hermitian manifolds. Guo-Phong [15] established the $L^{\infty}$-estimates for general classes of fully non-linear equations on Hermitian manifolds, their method can also be applied to open manifolds with a positive lower bound on their injectivity radii, the estimates on open manifolds rely on the behavior of the solution function $u$ in the interior of the open manifolds.
In this paper, we establish the $C^{0}$-estimates for (1.1) that depend only on the behavior of the solution function $u$ at infinity on the open manifolds. As we know, any Kähler form $\omega$ can be locally expressed by $\omega=\sqrt{-1}\partial\bar{\partial}\varphi>0$, where $\varphi$ is called the Kähler potential. The key idea of $C^{0}$-estimates is to assume that the complex manifold $M$ admits a bounded global Kähler potential, see the following definition:
@@
The following definition was introduced by Guan [13] and Székelyhidi [25], we restate it here for application in our case.
Definition 1.3.
We say that $u$ is a $\mathcal{C}$-subsolution for the equation $F(A)=h$ (see (1.1)) if the following holds. We require that for every point $x\in M$, if $\lambda=(\lambda_{1},\ldots,\lambda_{n})$ denote the eigenvalues of the endomorphism $\alpha^{i\bar{p}}g_{j\bar{p}}$ at $x$, then for all $i=1,\ldots,n$ we have
(1.7) $\lim_{t\rightarrow\infty}f(\lambda+t\textbf{e}_{i})>h(x)$
when $h(x)$ is independent of $u$. If $h(x,u)$ depends on both $x$ and $u$, we instead require
(1.8) $\lim_{t\rightarrow\infty}f(\lambda+t\textbf{{e}}_{i})=\infty.$
Cheng-Yau [4] derived the a priori estimates for complex Monge-Ampère equations on noncompact Kähler manifolds using generalized maximum principle. To solve the Gauduchon conjecture, Székelyhidi-Tosatti-Weinkovewe [24] established the second order estimates of a general class of fully non-linear equations on compact Hermitian manifolds. By combining the generalized maximum principle with certain results from [24], we obtain the a priori estimates for (1.1) up to $C^{k+\beta}$:
@@
Let $(M,\alpha)$ be a complete Hermitian manifold with bounded geometry (see Definition 3.1) of order k1k-1, $k\geq 4$. Fix a real $(1,1)$-form $\chi$ with bounded geometry of order k1k-1 on $(M,\alpha)$, given $h\in\tilde{C}^{k-2+\beta}(M)$, let $u\in\tilde{C}^{k+\beta}(M)$ satisfy that
@@
Let $\Omega$ be a complete Kähler manifold with a bounded global Kähler potential $\varphi<0$ and $\lim_{x\rightarrow\infty}\varphi(x)=0$. Let $\alpha$ be a Hermitian metric on $\Omega$ with bounded geometry of order k1k-1, where $k\geq 4$. Fix a real $(1,1)$-form $\chi$ with bounded geometry of order k1k-1, and $\chi$ is uniformly equivalent to $\alpha$ on $(\Omega,\alpha)$ (see section 2), given $h\in\tilde{C}^{k-2+\beta}(\Omega\times\mathbb{R})$, where $\beta\in(0,1)$. Consider the following equation on $\Omega$:
@@
Suppose that the above equation satisfies conditions (a1)-(a5) and $F(\alpha^{i\bar{p}}\chi_{j\bar{p}})=0$, assume that $|h(x,0)|=O(\underset{\min}{\lambda}(\alpha^{a\bar{l}}\varphi_{b\bar{l}}))$. Suppose 0 is a $\mathcal{C}$-subsolution of (1.10). Then the equation (1.10) exists a solution $u\in\tilde{C}^{k+\beta}(\Omega)$ such that $\chi+\sqrt{-1}\partial\bar{\partial}u$ is uniformly equivalent to $\alpha$.
If, in addition, assuming that $h_{u}(x,u)$ is strictly positive, we can generalize Theorem 1.5 to Hermitian manifolds:
@@
Suppose that the above equation satisfies conditions (a1)-(a5) and $F(\alpha^{i\bar{p}}\chi_{j\bar{p}})=0$, suppose 0 is a $\mathcal{C}$-subsolution of (1.11). Then the equation (1.11) exists a solution $u\in\tilde{C}^{k+\beta}(\Omega)$ such that $\chi+\sqrt{-1}\partial\bar{\partial}u$ is uniformly equivalent to $\alpha$.
@@
Theorem 1.7.
Let $\Omega$ be a strictly pseudoconvex domain in a complete Kähler manifold $(M,h)$ as in Proposition 3.5 with $l=\infty$. Define $g=-\log(-\varphi)$, where $\varphi$ is the defining function of $\Omega$, suppose that $|\partial\bar{\partial}\varphi|_{h}>c>0$. Fix a volume form $V(x)$ on $\Omega$ such that
(1.12) $|(\sqrt{-1})^{n}(V(x)-\det(g_{i\bar{j}}))dz^{1}\wedge d\bar{z}^{1}\cdots\wedge dz^{n}\wedge d\bar{z}^{n}|_{h}=O(\varphi^{2(1-n)})$
Then there exists a complete Kähler metric $\alpha$ on $\Omega$ such that $\det(\alpha_{i\bar{j}})=V(x)$.
@@
Denote the first Chern class of a Kähler manifold $M$ by $c_{1}(M)$. The existence of the Kähler-Einstein metrics on compact Kähler manifolds has been established in all cases: When $c_{1}(M)<0$, see Aubin [1] and Yau [36]; When $c_{1}(M)=0$, see Yau [36]; When $c_{1}(M)>0$, see Chen-Donaldson-Sun [6, 7, 8] and Tian [34, 35]. We obtain the following theorem on noncompact Kähler manifold:
Theorem 1.8.
Let $(M,\alpha)$ be a complete Kähler manifold with negative Ricci curvature bounded from below by a negative constant ($0>{\rm Ric}(\alpha)\geq c\alpha$ for some constant $c<0$). Then, for any compact subset $K\subset M$, there exists a Kähler-Einstein metric on $K$. Moreover, this metric can be extended to some open set $N\subset M$ and forms a complete Hermitian metric on $N$.
A related problem, the existence of Hesse-Einstein metrics ($\beta(g)=\lambda g$, see Section 8) on compact affine Kähler manifolds (Hessian manifolds) with negative first Chern class, was solved by Cheng-Yau [5]. Puechmorel-Tô [23] used geometric flow to provide another method for constructing the above metrics. For more studies on affine differential geometry, see ([3, 16, 22]). In the following, we use the a priori estimates for (1.1) to solve the real Monge-Ampère equations and construct the Hesse-Einstein metrics on certain noncompact affine manifolds.
Theorem 1.9.
Let $(\Omega,g)$ be a complete Hessian manifold with bounded geometry of infinity order. Suppose that the second Koszul form of $(\Omega,g)$ is bounded from below by a positive constant ($\kappa(g)>cg$ for some constant $c>0$), then $M$ admits a complete Hesse-Einstein metric.
The rest of this paper is organized as follows. In Section 2, we provide some preliminaries which may be used later. In Section 3, we construct some complete metrics with bounded geometry. In Section 4, we establish the $C^{0}$ estimates. In Section 5, we prove theorem 1.4. In Section 6, we solve the equation (1.1). In Section 7, we discuss the complex Monge-Ampère equations and some geometric applications. In Section 8, we consider some applications of (1.1) in affine geometry. We write an outline at the beginning of each section to introduce some key ideas.
@@ 2. Preliminaries
In this section, we provide some basic results and preliminaries which may be used in the following sections. Some notations and conventions used in this paper will also be introduced at various points throughout this section.
@@
$\bullet$ The tensor operations in this paper follow the Einstein summation convention.
$\bullet$ We extend covariant derivatives to act naturally on any type of tensor. For example, if $W$ is a tensor with components $W_{k}^{i\bar{j}}$, then define $\nabla_{m}W_{k}^{i\bar{j}}=\partial_{m}W_{k}^{i\bar{j}}+\Gamma_{lm}^{i}W_{k}^{l\bar{j}}-\Gamma_{km}^{l}W_{l}^{i\bar{j}}$ $\nabla_{\bar{m}}W_{k}^{i\bar{j}}=\partial_{\bar{m}}W_{k}^{i\bar{j}}+\overline{\Gamma_{lm}^{j}}W_{k}^{i\bar{l}}$
$\bullet$ The Hermitian metric $g$ defines a pointwise norm $|\cdot|_{g}$ on any tensor. For example, with $X,Y,a,b$ as above, we define $|X|_{g}^{2}=g_{i\bar{j}}X^{i}\overline{X^{j}},\;\;|Y|_{g}^{2}=g_{i\bar{j}}Y^{\bar{j}}\overline{Y^{\bar{i}}},\;\;|a|_{g}^{2}=g^{i\bar{j}}a_{i}\overline{a_{j}},\;\;|b|_{g}^{2}=g^{i\bar{j}}b_{\bar{j}}\overline{b_{\bar{i}}}.$ This is extended to any type of tensor. For example, with $W$ as above, we define $|W|^{2}_{g}=g^{k\bar{l}}g_{i\bar{j}}g_{p\bar{q}}W_{k}^{i\bar{q}}\overline{W_{l}^{j\bar{p}}}.$
$\bullet$ Noted that the Chern connection is a connection such that $\nabla g=\nabla J=0$. The associated Chern curvature tensor for the metric $g$ is then defined as $R_{i\bar{j}k}^{\;\;\;\;\;l}=-\partial_{\bar{j}}\Gamma_{ik}^{l}.$ The Chern-Ricci curvature is defined by $R_{i\bar{j}}=g^{k\bar{l}}R_{i\bar{j}k\bar{l}}=-\partial_{i}\partial_{\bar{j}}\log\det g.$ Note that if $g$ is not Kähler, then $R_{i\bar{j}}$ may not equal to $g^{k\bar{l}}R_{k\bar{l}i\bar{j}}$.
$\bullet$ For two real $(1,1)$-forms $\chi$ and $\chi^{\prime}$ on $(M,g)$, we say that $\chi$ is uniformly equivalent to $\chi^{\prime}$ if and only if $c^{-1}\chi^{\prime}<\chi<c\chi^{\prime}$ for some constant $c>0$.
$\bullet$ Suppose that $F$ is a function defined by $F(B)=f(\lambda(B))$ for a Hermitian matrix $B=\{B_{i\bar{j}}\}$ with $\lambda(B)\in\Gamma$, throughout the paper we shall use the notation $F^{ij}(B)=\frac{\partial F}{\partial B_{i\bar{j}}},\;\;\;\;\;F^{ij,kl}(B)=\frac{\partial^{2}F}{\partial B_{i\bar{j}}\partial B_{k\bar{l}}},$ where $B_{i\bar{j}}$ is a complex variable. If $f$ satisfies (a1) and (a2), then $F(B)$ is concave (see for instance Spruck [29] in the case of matrices with real entries). This means that: $F^{ij}(A)(B-A)_{i\bar{j}}\geq F(B)-F(A)\geq F^{ij}(B)(B-A)_{i\bar{j}},$ for any Hermitian matrices $A$ and $B$ such that $\lambda(A),\lambda(B)\in\Gamma$.
$\bullet$ We denote the minimum and maximum eigenvalues of $B$ by $\underset{\min}{\lambda}(B)$ and $\underset{\max}{\lambda}(B)$, respectively.
$\bullet$ For two real functions $f$ and $h$, we denote $|f|=O(|h|)$ if there exists a constant $C>0$ such that $|f|\leq C|h|$.
$\bullet$ For a complete noncompact Hermitian manifold $(M,g)$, we say $z\rightarrow\infty$ if and only if $d_{g}(z,z_{0})\rightarrow\infty$ for some fix point $z_{0}\in M$, where $d_{g}$ denotes the distance function induced by the metric $g$. Let $z_{1}\in M$ be another point, since $d_{g}(z,z_{1})\geq d_{g}(z,z_{0})-d_{g}(z_{1},z_{0})$, this definition is independent of the choice of $z_{0}$.